OpenStudy (anonymous):

Which method is most convenient in solving the given system of equations? x+3y=62x−3y=3

hero (hero):

x + 3y = 6 2x - 3y = 3

hero (hero):

@jlovely2, would elimination method or substitution method work best here?

OpenStudy (anonymous):

Yeah I would say substitution too. Solving the first equation for x in terms of y and z yields, x=3y-2z-12. Substituting the value of x in the second and third equations yields a system of two equations in two unknowns: (3y-2z-12)+2y+3z=6 2(3y-2z-12)-3y-z=-2 => 3y-2z-12+2y+3z=6 6y-4z-24-3y-z=-2 => 5y+z=18 3y-5z=22 for z gives, z=-5y+18 substituting the value of z in the second equation gives: 3y-5(-5y+18)=22 3y+25y-90=22 28y=112 y=4 substituting the value of y in the first equation: 5(4)+z=18 20+z=18 z=-2 substituting the value of y and z in the value of x that we first got ^^ : x=3y-2z-12 x=3(4)-2(-2)-12 x=12+4-12 x=4 The solution set is: { (4,-2,4) }

OpenStudy (anonymous):

did i help u

hero (hero):

@jonjenkins7653, you solved it good and well, but you used the wrong method and that wasn't what the question asked for.